Circular No. 8063

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003aa IN NGC 3367 Further to IAUC 8061, B. Swift, H. Pugh, and W. Li report the LO-

TOSS discovery of an apparent supernova (mag ∼ 17.6) on unfiltered KAIT images taken on Jan. 31.4 and Feb. 1.4 UT. SN 2003aa is located at α = 10^{h}46^{m}36^{s}.82, δ = +13^{o}45′32′′.2 (equinox 2000.0), which is 26′′.4 east and 30′′.2 north of the nucleus of NGC 3367. A KAIT image taken on Jan. 24.4 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVAE 2003Y AND 2003Z T. Matheson, P. Challis, and R. Kirshner, Center for Astrophysics,

report that a spectrum (range 370–750 nm) of SN 2003Y (cf. IAUC 8062), obtained by M. Calkins on Jan. 31.34 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Several titanium features appear stronger than normal, indicating that this may be a subluminous event. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 5079 km/s for the host galaxy, the supernova expansion velocity is 10400 km/s for Si II (rest 635.5 nm). A spectrogram of SN 2003Z (cf. IAUC 8062), obtained on Jan. 31.36, shows it to be a type-II supernova. The spectrum consists of a very blue continuum with well-defined P-Cyg lines of hydrogen and helium. Adopting the NED recession velocity of 1289 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6800 km/s.

IGR J16318–4848 T. J.-L. Courvoisier, INTEGRAL Science Data Centre (ISDC), Ver-

soix, writes that R. Walter, ISDC; J. Rodriguez, Centre d’Etude Atomique and Service d’Astrophysique, Saclay; L. Bouchet, Centre d’Etude Spatial du Rayonnement, Toulouse; and A. A. Lutovinov, Space Research Institute, Russian Academy of Sciences, Moscow, report the detection of the first transient source observed by the ISGRI detector of the IBIS instrument onboard the European Space Agency’s INTEGRAL satellite. The source appeared at α = 16^{h}31^{m}52^{s}, δ = −48^{o}48′.5 (equinox 2000.0; uncertainty ∼ 2′), while INTEGRAL was performing an observation of the Galactic plane on Jan. 29.271 UT. The source flux in the band 15–40 keV was 50–100 mCrab. The source was observed to vary significantly on a timescale of 1000 s. INTEGRAL will observe the source again during Feb. 1–3. Observations at other wavelengths are encouraged.

2003 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green